OpenStudy (iwanttogotostanford):

HELP ASAP 2 QUESTIONS

OpenStudy (iwanttogotostanford):

@sooobored @Austin.L

OpenStudy (iwanttogotostanford):

@Nnesha @Directrix

OpenStudy (sooobored):

if given \[f(x)=x^{n}\] the derivative \[\frac{df(x)}{dx}=f'(x)=nx^{n-1}\]

OpenStudy (iwanttogotostanford):

so the first would be 8?

OpenStudy (sooobored):

just by looking and doing it roughly in my head, i dont think so

OpenStudy (iwanttogotostanford):

what do you think....

OpenStudy (sooobored):

\[f(x)=1/x\] \[f'(x)=-1*x^{-2}\] if x = -1 then f'(x) =-1

OpenStudy (sooobored):

it helps to write down the steps

OpenStudy (sooobored):

sorry, forgot to mention \[f(x)=1/x= x^{-1}\]

OpenStudy (iwanttogotostanford):

ah ok what do you think the solutions are to the 2 would be?

OpenStudy (sooobored):

well, you can still apply the derivative power rule that i wrote above for number 2

OpenStudy (sooobored):

some other derivative rules, if you dont remember them given a function f(x)=c*x^n the derivative \[f'(x)=c*nx^{n-1}\] also if f(x)=g(x)+h(x) then the derivative \[f'(x)=g'(x)+h'(x)\]

OpenStudy (iwanttogotostanford):

can you just help me find the answer thats how i learn the best honestly... :(

OpenStudy (sooobored):

ill help you break it up into simpler problems what is the derivative of -12x^2 ?

OpenStudy (iwanttogotostanford):

i just thing -8 is the answer to 1

OpenStudy (iwanttogotostanford):

@mathmate

OpenStudy (iwanttogotostanford):

@retirEEd

OpenStudy (iwanttogotostanford):

@3mar need help so glad you're here please help!

OpenStudy (retireed):

I think I did this yesterday for you.

OpenStudy (3mar):

Well, I am here.

OpenStudy (3mar):

For the 1st one : it is -8 CORRECT!

OpenStudy (3mar):

And for the 2nd: It will be: \[f(x)=-12x^2+9x\] \[\frac{ df(x) }{ dx }=\frac{ d }{ dx }(-12x^2)+\frac{ d }{ dx }(9x)\] \[\frac{ df(x) }{ dx }=(-12)(2)(x^{2-1})+(9)(x^{1-1})=-24x+9\] then you can put x=6: \[f'(x)=-24(6)+9=-135\] Sorry for being late! Hope that helps

OpenStudy (3mar):

Please, tell me that the answer reaches you and you get satisfied! Salam!